OpenStudy (anonymous):

how do you solve [(a-b)/(c-d)]^2

OpenStudy (campbell_st):

well it would need to be equal to something to be able to solve it

TheSmartOne (thesmartone):

\[[\frac{ (a-b) }{ (c-d) }]^2=\frac{ (a-b)^2 }{(c-d)^2 }=\frac{ a^2-2ab+b^2 }{ c^2-2cd+d^2 }\]